Show me how to solve this problem
Multiply in the indicated base
12[sub]5[/sub]x 4[sub]5[/sub]
1
Expert's answer
2011-05-24T07:16:06-0400
Let's change the base to 10 and multiply the numbers in usual way:
125 = 5*1 + 2 = 7
45 = 4
7x4 = 28.
2810 = 5*5 + 3 = 535
Answer: 535.